A. The pharmacist receives the following prescription. Rx: 0.5% 10 mL Tetracaine Hydrochloride 0.2% Epinephrine Bitartrate in NSS Sodium Chloride Purified Water ad Make isotonic Solution 15 mL Normal Saline Solution How many milligrams of sodium chloride are needed to make this solution isotonic?
Answer #1

1. a)100ml solution has 0.5gm of tetracaine hydrochloride.

Then 15ml solution has 0.5/100×15= 0.075gm of tetracaine hydrochloride.

b) E- value of tetracaine hydrochloride = 0.18

0.075gm ×0.18=0.0135gm of NaCl is represented by tetracaine hydrochloride.

Volume of solution :-

15ml (total value)- 10ml (volume of epinephrine bitartrate) =5ml.

0.9gm of NaCl is present in 100ml of solution

Then amount of NaCl in 5ml = 0.9÷100 ×5= 0.045gm of NaCl is present in 5ml of isotonic solution.

The actual amount of NaCl is 0.045 gm - 0.0135gm =0.0315gm of NaCl only.

0.9gm of NaCl is present in 100ml

So, 0.0315gm of NaCl is present in 0.9÷100 × 0.0315 = 3.5 ml

3.5 ml of Sodium chloride is required to make this isotonic solution .